OpenStudy (anonymous):

Eighty-five more than the square of a number is the same as the square of the quantity that is 17 less than the number. What is the number? Eighty-five more than the square of a number is the same as the square of the quantity that is 17 less than the number. What is the number? @Mathematics

OpenStudy (anonymous):

If I follow it correctly: x^2 + 85 = (x-17)^2 85 = -34x + 289 x = 6